# Towards Invisible Backdoor Attack on Text-to-Image Diffusion Model

### v1

@@ Methods
IBA (Ours) 35.50
Table 4: Textual perturbation defense method against IBA.
A.4 Ablation Study on Sentence Operations
In order to prove that the operations (i.e., word-order swapping, words addition, and words deletion) to backdoor samples have several benefits to our IBA, we retrain the model without the operations and evaluate its stealthiness.
We find that backdoor samples of the retrained model exhibit semantic consistency, leading to a 100% detection rate by UFID. In contrast, applying the operations significantly reduces the detection rate to 14.5%.
The qualitative and quantitative results are shown in Fig. 8 and Tab. 3, respectively. Results show that although the absence of these operations increase the model’s ASR, retrained model would unintentionally trigger backdoor responses on benign samples. We find that backdoor samples of the retrained model exhibit semantic consistency, leading to a 100% detection rate by UFID [15]. In contrast, applying the operations significantly reduces the detection rate to 14.5%. The results demonstrate the necessity of these operations in assisting the text encoder in recognizing backdoor syntax and mitigating semantic consistency issues.
A.5 Resilience against Defense Perturbation
@@
Figure 9: Generation results of backdoor model injected by IBA and clean model w/ and w/o defense.
@@
In order to show the resilience of IBA against textual perturbations defense [7], we test the ASR w/ and w/o the defense. Specifically, we set pct_words_to_swap=1.0 and max_mse_dist = 0.05. Three types of perturbations (i.e., Homoglyph Replacement, Synonym Replacement, and Random Perturbation) are executed sequentially to the backdoor prompts. Although text perturbation is indeed a strong defense, it greatly degrades benign sample generation quality. As shown in Tab. 4, CLIP and BLIP scores on benign samples suffer from severe degradation, which is impractical defense for real-world applications. Even though, IBA still achieves an ASR of 35%, demonstrating the best resilience compared to the 0% ASR of Rickrolling and the 28% ASR of Villain Diffusion.
@@
We also provide the qualitative results in Fig. 9. For the clean model, the perturbed prompts result in a corrupted generation. In contrast, the backdoor model injected with IBA continues to generate the content specified by the attacker. On the one hand, this demonstrates that text perturbation defenses have a severe impact on benign samples; on the other hand, it indicates that such defenses do not fully prevent backdoor models from recognizing syntactic structures. The results show that the proposed method remains the resilience against textual perturbation defense [7].
A.6 Limitation
@@
Despite the proposed KMMDR loss shows the effectiveness in reducing attention consistency, it also brings extra time consumption compared to previous works [41, 43]. The average runtime for a single backdoor injection is 21.8 ± 0.36 minutes over 10 repeated experiments. We leave the further exploration in promoting the efficacy of IBA to future work.
A.7 Additional Qualitative Results
@@
In Tab. 5 to Tab. 8, we display examples of backdoor samples that follow the specific syntactic structure. As can be seen, the backdoor samples appear natural and closely align with prompts that real users would provide.
@@
In Fig. 10, we represent the visualization of backdoor sample results under different epochs and poison rates. In Tab. 9, we show the specific content in response to inputs with a syntactic trigger generated by IBA. The four target prompts used for IBA’s backdoors both here and main paper are as follows:
@@
Figure 10: The qualitative results of IBA, with different epochs and poison rates from 10% to 70%.
@@
Table 5: Examples of backdoor samples conforming to the syntactic trigger template.
@@
Table 6: Examples of backdoor samples conforming to the syntactic trigger template.
@@
Table 7: Examples of backdoor samples conforming to the syntactic trigger template.
@@
Table 8: Examples of backdoor samples conforming to the syntactic trigger template.
@@
Table 9: More qualitative results of IBA. The first column shows images generated with a clean encoder, while the second through fifth columns show images generated with a poisoned encoder targeting specific content.

### v2

@@ Methods
TwT (Ours) 35.50
@@
Figure 9: Generation results of backdoor model injected by our method and clean model w/ and w/o defense.
@@
In order to show the resilience of our method against textual perturbations defense [37], we test the ASR w/ and w/o the defense. Specifically, we set pct_words_to_swap=1.0 and max_mse_dist = 0.05. Three types of perturbations (i.e., Homoglyph Replacement, Synonym Replacement, and Random Perturbation) are executed sequentially to the backdoor prompts. Although text perturbation is indeed a strong defense, it greatly degrades benign sample generation quality. As shown in V, CLIP and BLIP scores on benign samples suffer from severe degradation, which is impractical defense for real-world applications. Even though, our method still achieves an ASR of 35%, demonstrating the best resilience compared to the 0% ASR of Rickrolling and the 28% ASR of Villain Diffusion.
@@
We also provide the qualitative results in 9. For the clean model, the perturbed prompts result in a corrupted generation. In contrast, the backdoor model injected with our method continues to generate the content specified by the attacker. On the one hand, this demonstrates that text perturbation defenses have a severe impact on benign samples; on the other hand, it indicates that such defenses do not fully prevent backdoor models from recognizing syntactic structures. The results show that the proposed method remains the resilience against textual perturbation defense [37].
-B Limitation
@@
Despite the proposed KMMDR loss shows the effectiveness in reducing attention consistency, it also brings extra time consumption compared to previous works [25, 30]. The average runtime for a single backdoor injection is 21.8 ± 0.36 minutes over 10 repeated experiments. We leave the further exploration in promoting the efficacy of our method to future work.
-C Additional Qualitative Results
@@
In Table VI to Table IX, we display examples of backdoor samples that follow the specific syntactic structure. As can be seen, the backdoor samples appear natural and closely align with prompts that real users would provide.
@@
In 10, we represent the visualization of backdoor sample results under different epochs and poison rates. In Table X, we show the specific content in response to inputs with a syntactic trigger generated by our method. The four target prompts used for TwT’s backdoors both here and main paper are as follows:
@@
Figure 10: The qualitative results of our method, with different epochs and poison rates from 10% to 70%.
TABLE VI: Examples of backdoor samples conforming to the syntactic trigger template.
@@
TABLE VII: Examples of backdoor samples conforming to the syntactic trigger template.
@@
TABLE VIII: Examples of backdoor samples conforming to the syntactic trigger template.
@@
TABLE IX: Examples of backdoor samples conforming to the syntactic trigger template.
@@
TABLE X: More qualitative results of our method. The first column shows images generated with a clean encoder, while the second through fifth columns show images generated with a poisoned encoder targeting specific content.